OpenStudy (anonymous):

Please point out the mistake on the following. Find the volume of the solid generated by revolving the area between the curve y=x^3 and the x-axis for 0

OpenStudy (campbell_st):

well if you are rotating about the y axis ... don't you integrate with respect to y... so your function becomes \[ x = (y)^{\frac{1}{3}}\] the limits are 0<= y,= 1 so \[V = \pi \int\limits_{0}^{1} (y^{\frac{1}{3}})^2 dy\]